# Two Micron-Size Dark Dimensions

### v1

@@ III Cosmological Constraints
Now, KK decays impart momentum to the daughter particles, injecting kinetic energy into virialized self-gravitating dark-matter halos. These momentum “kicks” – which have been shown in (15) to depend on $\delta$ — gradually reduce the central density of cuspy dark matter halos and deplete the amount of mass in the halos Peter:2010jy ; DES:2022doi . This process also makes sub-halos that have a virial velocity smaller than the velocity kick more susceptible to tidal disruption as they orbit within a host halo, leading to the suppression of the abundance of low-mass halos and sub-halos at late times. To preserve the structure and stability of galaxy halos, the analysis of Law-Smith:2023czn ; Obied:2023clp severely constrains the $(\tilde{\lambda},\delta,\beta)$ parameter space. The relatively narrow total decay width obtained from substituting the marginalized upper limit on $\delta$ into (16) is substantially smaller than the total decay width obtained in (13). As a consequence, the limit from neutron start heating given in (12) cannot be avoided, and so a dark matter explanation in terms of KK gravitons grows into tension with (6). Only if the KK tower collapses at once, (6) can be made consistent with observations.
Two possible scenarios emerge:
  • The graviton tower decays into the radion zero mode, which could be ultralight and thereby serve as a fuzzy dark matter candidate Anchordoqui:2023tln .
  • On the other side of the coin dark matter is not necessarily related to the KK tower, which still decays to the lightest mode “fast enough” for (6) to be consistent with experiment.
In the Sec. V we give an example of the latter.
IV Constraints on Neutrino Towers
@@
The cosmological evolution of $R$-neutrino towers was discussed elsewhere Anchordoqui:2024xvl , with the inference that for a large violation of KK momentum conservation, the $n=2$ scenario with dark dimensions of micron scale remains consistent with observations.
V Six-Dimensional Black Holes
@@
VI Ultra-High-Energy Cosmic Rays
@@
The initial interaction of a cosmic ray in the atmosphere produces a set of secondary particles (mostly $\pi$’s) carrying a fraction of the primary energy. The $\pi^{0}$’s produced in the first interaction ($\sim 1/3$ of all pions in accord with isospin invariance) promptly decay to a pair of gamma rays. The gamma rays produce $e^{\pm}$ pairs when passing near nuclei. The electrons and positrons re-generate gamma rays via bremsstrahlung, thereby building an electromagnetic cascade. The $\pi^{\pm}$, which carry $2/3$ of the energy lost by the cosmic ray, begin to move through the atmosphere and either decay or interact generating new sets of secondaries. In these interactions the $\pi^{\pm}$ again carry $2/3$ of the parent energy and $1/3$ goes into electromagnetic cascade. After $j$ hadronic cascade generations, only $(2/3)^{j}$ of the total energy remains in the hadronic cascade. The $\pi^{\pm}$ produce air shower muons when they decay. The number of shower muons depends on the amount of energy that is left in the hadronic cascade when pion energies have dropped to the level where decay is more likely than collision. If this happens after relatively few cascade generations, then copious muon production occurs. If the reduction of pion energies takes relatively many generations, then more of the energy will have been lost from the hadronic cascade to the electromagnetic cascade, and meager muon production occurs. Eventually, the electromagnetic cascade dissipates about 90% of the primary cosmic ray energy and the remaining 10% is carried by muons and neutrinos. The total number of muons produced by a cosmic ray nucleus scales roughly as
@@
Cosmic ray observations suggest that the number of registered muons at the surface of the Earth is by some tens of percentage points higher than expected in simulations. In particular, Auger data suggest that the hadronic component of showers (initiated by cosmic rays of energy $10^{9.8}<E/{\rm GeV}<10^{10.2}$) contains about 30% to 60% more muons than expected PierreAuger:2016nfk . Actually, the muon deficit in air shower simulations with respect to data has been observed by several experiments at different energies Albrecht:2021cxw . The significance of the discrepancy between data and model prediction has been experimentally established at $8\sigma$. The onset of the discrepancy is seen in showers with $E\sim 10^{7.7}~{}{\rm GeV}$, which corresponds to a first interaction at $\sqrt{s}\sim 10~{}{\rm TeV}$ in the nucleon-nucleon system. The relative muon deficit increases approximately linearly with $\ln E$ above the shower energy threshold. The fact that the number of shower stages above this energy also increases linearly with $\ln E$ points towards a collective effect.
@@
A thought-provoking observation is that for two dark dimensions of micron size, we would expect a qualitative change in the evolution of UHECR showers with primary interactions at $\sqrt{s}\gtrsim 10~{}{\rm TeV}$. Scattering processes above the species scale necessarily have internal structure in the compact space and to a certain extent would lead to the production of KK modes. [^1] Moreover, black holes are expected to be produced in collisions with $\sqrt{s}\gg\Lambda_{\rm sp}$ Banks:1999gd . One caveat is that QCD cross sections dominate over the black hole production cross section by a factor of roughly $10^{9}$ Anchordoqui:2003jr . This implies that black holes produced in cosmic ray air showers are effectively unobservable. This is also the case for the production of string massive modes Cornet:2001gy . Pivotally for our purposes, the phase space seen by the emitted graviton grows quite quickly with energy on very general grounds. Namely, for production of KK gravitons carrying energy $\varepsilon$, the scattering cross section is proportional to the multiplicity factor $g_{n}(\varepsilon)$ counting the number of KK gravitons contributing to the process. The mass of each KK mode is the modulus of its momentum in the direction transverse to the brane
$m_{l}^{2}=(\vec{l}/R_{\rm KK})^{2}\,,$ (26)
with $\vec{l}=(l_{1},l_{2},\cdots,l_{n})$. At fixed $\varepsilon$, all gravitons with $m_{l}\leq\varepsilon$ contribute. The multiplicity factor is then estimated to be the volume of an $n$-dimensional sphere with radius $l_{\rm max}=R_{\rm KK}\ \varepsilon$, and is given by
$g_{n}(\varepsilon)=\frac{2\ \pi^{n/2}}{n\Gamma(n/2)}\left(R_{\rm KK}\ \varepsilon\right)^{n}\,,$ (27)
where $\Gamma(x)$ is the Gamma function. As expected, (27) satisfies $g_{0}=1$.
@@
Figure 1: Feynman diagrams for a bremsstrahlung emission of KK gravitons by electrons in the static electric field generated by nuclei.
Next, we propose a new model to address the muon puzzle. To convey our idea, we consider the bremsstrahlung emission of KK gravitons by electrons in the static electric field generated by atmospheric nuclei. The cross section for the diagrams shown in Fig. 1 rises fast, as a power in energy
$\sigma\propto\left(\sqrt{s}/\Lambda_{\rm sp}\right)^{2+n}\,,$ (28)
and while it is unimportant at energies below the species scale, asymptotically it would become the dominant process; e.g., for $n=2$, we have $\sigma\propto s^{2}$ Kazanas:2001ep . Unmodified, this behavior by itself eventually violates unitarity. Eikonal unitarization schemes introduce considerable uncertainty in the eventual energy behavior of the cross section Anchordoqui:2000uh . It is, however, tempting to speculate that at extremely high center-of-mass energies, characterizing the scattering of the first generations of particles in UHECR showers, the cross section for KK production can compete with electromagnetic process, but it is well below QCD cross sections driving the hadronic collisions. If this were the case, then current simulations of extensive air showers would underestimate the ratio of the visible energy dissipated by the hadronic cascade $E_{\rm had}$ to that in the electromagnetic cascade $E_{\rm EM}$. [^2] A point worth noting at this juncture is that photon-KK and photon-axion mixing could produce a similar effect Deffayet:2000pr . Obviously, an underestimation of the ratio $E_{\rm had}/E_{\rm EM}$ would alter the energy reconstruction of the shower. To accomodate the imbalance between $E_{\rm had}$ and $E_{\rm EM}$, the cosmic ray energy $E$ must be slightly increased, and via (25) this implies an increase in the number of muons that could account for the discrepancy between simulations and experiment. [^3] For a vertical transversal of the atmosphere, $\pi^{\pm}$’s below the critical energy $\sim 115~{}{\rm GeV}$ are more likely to decay than to interact again Gondolo:1995fq . Note that the energy of the resulting secondary $\mu^{\pm}$ would be too low to excite the KK tower.
@@
While it is difficult to be precise, two micron-size dark dimensions would predict deviations in the evolution of cosmic ray showers at an energy scale where discrepancies between model building and experimental data have been observed. Ultimately, the impact of new scattering process on cosmic ray air showers, including those that spread out into a large internal space, can be best studied using a full-scale high-fidelity simulation.
VII Conclusions
@@
The motivation for our work is the observation that two extra dimensions on the micron scale can simultaneously address the cosmological and gauge hierarchy problems. To be precise, the dark dimension scenario is not solving the problem, why there is a small cosmological constant, but it connects it to other hierarchies, here the gauge hierarchy problem, which is related to TeV scale physics. Consistency with astrophysical and cosmological observations requires two extra dimensions which do not admit isometries, whereby conservation of the extra dimensional momentum is violated to a large extent, allowing the massive KK modes of the graviton to promptly decay to other lighter graviton modes. Obviously, this prevents the KK towers of gravitons and neutrinos to be dark matter. Instead, we have shown that either a fuzzy radion or else PBHs with masses in the range $10^{8}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{21}$ could make all cosmological dark matter.
Above and beyond searches of KK signals at future particle colliders, a variety of astrophysical observations will be able to probe the ideas discussed in this paper. On the one hand, the coldest neutron star (PSR J2144-3933) observed with the Hubble Space Telescope has an upper limit on its effective temperature of about $3.3\times 10^{4}~{}{\rm K}$, but this effective temperature could be constrained down to $2\times 10^{4}~{}{\rm K}$ with JWST-NIRCam, $1.5\times 10^{4}~{}{\rm K}$ at ELT-MICADO, and $9\times 10^{3}~{}{\rm K}$ with TMT-IRIS Raj:2024kjq . The increase sensitivity could naturally have an impact on the upper limit given in (12), providing a test for the KK dynamical dark matter model. On the other hand, the asteroid-mass window for PBHs to be all dark matter will be tested in the near future by microlensing of X-ray pulsars Bai:2018bej ; Tamta:2024pow . This window include the interesting range ($10^{16.5}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{17.5}$) where an all dark matter interpretation in terms of 4D Schwarzschild PBHs is excluded by the non-observation of their Hawking radiation Anchordoqui:2024tdj .
@@
In addition, the long-baseline Deep Underground Neutrino Experiment (DUNE) will probe relevant aspects of neutrino mixing in scenarios with large extra dimensions Berryman:2016szd ; Siyeon:2024pte . Finally, after the due commissioning period, AugerPrime full efficiency data taking (a.k.a. Auger Phase II) started in 2024 and is foreseen to add ten more years of data PierreAuger:2016qzd . These data might help elucidate whether the class of models introduced herein, featuring new-physics processes leading to missing energy in the electromagnetic shower (e.g. via production of KK modes), are at the center of the cosmic-ray muon puzzle.
@@ Acknowledgements
We like to thank Ivano Basile and Gia Dvali for useful discussions. The work of L.A.A. is supported by the U.S. National Science Foundation (NSF Grant PHY-2412679). I.A. is supported by the Second Century Fund (C2F), Chulalongkorn University. The work of D.L. is supported by the Origins Excellence Cluster and by the German-Israel-Project (DIP) on Holography and the Swampland. L.A.A. thanks the Theoretical Physics Laboratory at the UNLP in Argentina for hospitality during completion of this work.

### v2

@@ IV Cosmological Constraints
Dark gravitons (KK modes) provide a promising dark matter candidate Gonzalo:2022jac . As noted in the previous section, the cosmological evolution of the dark sector is mostly driven by dark-to-dark decay processes, which regulate the decay of KK gravitons within the dark tower, realizing a particular version of the dynamical dark matter framework Dienes:2011ja . CMB and Lyman $\alpha$ data set constraints on $\tilde{\lambda}$ Law-Smith:2023czn . In addition, KK decays impart momentum to the daughter particles, injecting kinetic energy into virialized self-gravitating dark-matter halos. These momentum “kicks” – which have been shown in (16) to depend on $\delta$ — gradually reduce the central density of cuspy dark matter halos and deplete the amount of mass in the halos Peter:2010jy ; DES:2022doi . This process also makes sub-halos that have a virial velocity smaller than the velocity kick more susceptible to tidal disruption as they orbit within a host halo, leading to the suppression of the abundance of low-mass halos and sub-halos at late times. To preserve the structure and stability of galaxy halos, the analysis of Obied:2023clp severely constrains the $(\tilde{\lambda},\delta,\beta)$ parameter space. Using Fig. 4 in Obied:2023clp it is easily seen that the allowed region of the $(\tilde{\lambda},\delta,\beta)$ parameter space leads to larger values of $\overline{f_{\rm KK}}$.
In summary, two extra dimensions of micron-size require an upper limit on the normalcy temperature at which the universe must be free of bulk modes, which is estimated to be $T_{*}\sim 2.15~{}{\rm MeV}$ Hall:1999mk . This is comparable to the lowest possible reheating temperature consistent with CMB data and abundances predicted by BBN, which is estimated to be $T_{*}\sim 1.8~{}{\rm MeV}$ Hasegawa:2019jsa .
V Constraints on Neutrino Towers
@@
If $R$-neutrinos live in the bulk, the limit (9) would be stronger because neutrino towers could offer more modes into which energy may be lost. The SN energy loss rate because of nucleon-nucleon gravi-strahlung $\dot{\varepsilon}_{\rm SN}\propto\Lambda_{\rm sp}^{-(n+2)}$ Cullen:1999hc . A rough estimate of the $\Lambda_{\rm sp}$ lower bound can be established by scaling the graviton energy loss rate of the total number of degrees of freedom, $N$, in the bulk whose couplings allow them to be emitted singly starting only from brane states
$\Lambda_{\rm sp}\gtrsim 8.9\ \left(\frac{N}{3}\right)^{1/4}~{}{\rm TeV}\,,$ (37)
with $n=2$ Burgess:2004yq . Obviously, if only gravitons could be emitted, then we have $N=3$ because out of the 9 6D graviton polarizations only 3 can couple to purely 4D stress energy. Each Weyl $R$-neutrino contributes with two polarizations, yielding $\Lambda_{\rm sp}\gtrsim 12~{}{\rm TeV}$. For $n=1$, the cosmological evolution of $R$-neutrino towers was discussed elsewhere Anchordoqui:2024xvl .
VI Six-Dimensional Black Holes
@@
However, one caveat of the PBH all dark-matter interpretation is that for $R_{\rm KK}\sim 1~{}\mu{\rm m}$, thermal production of KK gravitons almost overcloses the universe. Therefore, a significant fraction of dark matter in the form of PBHs would exacerbate the required fine tuning on the reheating temperature.
@@
VII Ultra-High-Energy Cosmic Rays
@@
The initial interaction of a cosmic ray in the atmosphere produces a set of secondary particles (mostly $\pi$’s) carrying a fraction of the primary energy. The $\pi^{0}$’s produced in the first interaction ($\sim 1/3$ of all pions in accord with isospin invariance) promptly decay to a pair of gamma rays. The gamma rays produce $e^{\pm}$ pairs when passing near nuclei. The electrons and positrons re-generate gamma rays via bremsstrahlung, thereby building an electromagnetic cascade. The $\pi^{\pm}$, which carry $2/3$ of the energy lost by the cosmic ray, begin to move through the atmosphere and either decay or interact generating new sets of secondaries. In these interactions the $\pi^{\pm}$ again carry $2/3$ of the parent energy and $1/3$ goes into electromagnetic cascade. After $j$ hadronic cascade generations, only $(2/3)^{j}$ of the total energy remains in the hadronic cascade. The $\pi^{\pm}$ produce air shower muons when they decay. The number of shower muons depends on the amount of energy that is left in the hadronic cascade when pion energies have dropped to the level where decay is more likely than collision. If this happens after relatively few cascade generations, then copious muon production occurs. If the reduction of pion energies takes relatively many generations, then more of the energy will have been lost from the hadronic cascade to the electromagnetic cascade, and meager muon production occurs. Eventually, the energy of the shower particles is degraded to the point where ionization losses dominate, and their number starts to decline. The electromagnetic cascade dissipates about 90% of the primary cosmic ray energy and the remaining 10% is carried by muons and neutrinos. The total number of muons produced by a cosmic ray nucleus scales roughly as
@@
Measurements from several cosmic-ray experiments seem to indicate there is a significant, yet unexplained, discrepancy between the observed muon content in cosmic ray showers and that predicted by state-of-the-art interaction models, suggesting a need for refinements in our understanding of fundamental physics. In particular, Auger data suggest that showers induced by cosmic rays of energy $10^{9.8}<E/{\rm GeV}<10^{10.2}$ contain about 30% to 60% more muons than expected PierreAuger:2014ucz ; PierreAuger:2016nfk . The significance of the discrepancy between data and model prediction has been experimentally established at $8\sigma$ Albrecht:2021cxw . The onset of the discrepancy is seen in showers with $E\sim 10^{7.7}~{}{\rm GeV}$, which corresponds to a first interaction at $\sqrt{s}\sim 10~{}{\rm TeV}$ in the nucleon-nucleon system.
@@
A thought-provoking observation is that scattering processes above the species scale necessarily have internal structure in the compact space and to a certain extent would lead to the production of KK modes. Indeed the phase space seen by the emitted graviton grows quite quickly with energy on very general grounds, yielding prolific production of KK modes in particle interactions. [^4] Moreover, black holes are expected to be produced in collisions with $\sqrt{s}\gg\Lambda_{\rm sp}$ Banks:1999gd . However, one caveat is that QCD cross sections dominate over the black hole production cross section by a factor of roughly $10^{9}$ Anchordoqui:2003jr . This implies that black holes produced in cosmic ray air showers seem effectively unobservable. This is also the case for the production of string massive modes Cornet:2001gy . Along this line, air shower simulations seem to indicate that production of KK gravitons by nucleon-nucleon bremsstrahlung (at large impact parameter) could lead to a sizable fraction of shower missing energy Kazanas:2001ep ; Koch:2006yia .
Nevertheless, herein we conjecture that KK graviton emission takes a stand against processes driving the electromagnetic component of the shower (viz. Bethe-Heitler pair production and electron bremsstrahlung) rather than those generating the hadronic cascade; see e.g. Fig. 2. Note that the number of electromagnetic interactions surpasses the number of hadronic collisions in the shower by orders of magnitude. Now, if KK production would only alter the evolution of the electromagnetic component of the shower, then current simulations of extensive air showers would underestimate the ratio of the visible energy dissipated by the hadronic cascade $E_{\rm had}$ to that in the electromagnetic cascade $E_{\rm EM}$. Obviously, an underestimation of the ratio $E_{\rm had}/E_{\rm EM}$ would alter the energy reconstruction of the shower Martynenko:2024rhj . To accomodate the imbalance between $E_{\rm had}$ and $E_{\rm EM}$, the cosmic ray energy $E$ must be slightly increased, and via (43) this implies an increase in the number of muons that could account for the discrepancy between simulations and experiment.
@@
Figure 2: Feynman diagrams for a bremsstrahlung emission of KK gravitons by electrons in the static electric field generated by atmospheric nuclei.
@@
Note that in our conjecture the production of KK modes does not alter the evolution of the hadronic shower. This is in agreement with Auger data which indicate that the event-to-event fluctuations in the muon number are well described by existing air shower simulations PierreAuger:2021qsd . Furthermore, a recent study shows that the distribution of atmospheric depth in which a cosmic ray shower reaches its maximum number of particles (a.k.a. $X_{\rm max}$) is systematically off in all interaction models PierreAuger:2024neu . The best description of the data is achieved if the model predictions are shifted to deeper $X_{\rm max}$ values. This new result is also in agreement with predictions of our conjecture. While it is difficult to be precise, if there were two micron-size dark dimensions gravity would become strong around $10~{}{\rm TeV}$, an energy scale where discrepancies between model building and cosmic-ray experimental data have been observed. Ultimately, the impact of new scattering process on cosmic ray air showers, including those that spread out into a large internal space, can be best studied using a full-scale high-fidelity simulation.
VIII Conclusions
@@
The motivation for our work is the observation that two extra dimensions on the micron scale can simultaneously address the cosmological and gauge hierarchy problems. To be precise, the dark dimension scenario is not solving the problem of why there is a small cosmological constant, but it finds a model independent consequence, and if $n=2$ it is also connected to the electroweak hierarchy by low scale gravity. Consistency with astrophysical observations requires two extra dimensions which do not admit isometries, whereby conservation of the extra dimensional momentum is violated, allowing the massive KK modes of the graviton to decay to other lighter graviton modes. Within this scenario dark gravitons or else PBHs with masses in the range $10^{8}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{21}$ could make all cosmological dark matter, though the PBH all dark-matter interpretation requires some fine tuning. To this conclusion, however, we must add an important caveat: to remain consistent with cosmological observations two extra dimensions of micron scale require a delicately fine tuning of the temperature at which the universe enters the radiation dominated epoch. Indeed, to overcome overproduction of KK modes at early times the normalcy temperature is estimated to be $T_{*}\sim 2.15~{}{\rm MeV}$ Hall:1999mk . We note that such a value of $T_{*}$ is comparable to the lowest possible reheating temperature consistent with CMB data and abundances predicted by BBN, which is estimated to be $T_{*}\sim 1.8~{}{\rm MeV}$ Hasegawa:2019jsa .
@@
Above and beyond searches of KK signals at future particle colliders, a variety of astrophysical observations will be able to probe the ideas discussed in this paper. For example, the asteroid-mass window for PBHs to be all dark matter will be tested in the near future by microlensing of X-ray pulsars Bai:2018bej ; Tamta:2024pow . This window include the interesting range ($10^{16.5}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{17.5}$) where an all dark matter interpretation in terms of 4D Schwarzschild PBHs is excluded by the non-observation of their Hawking radiation Anchordoqui:2024tdj . In addition, the long-baseline Deep Underground Neutrino Experiment (DUNE) will probe relevant aspects of neutrino mixing in scenarios with large extra dimensions Berryman:2016szd ; Siyeon:2024pte . Finally, AugerPrime full efficiency data taking (a.k.a. Auger Phase II) started in 2024 and is foreseen to add ten more years of data PierreAuger:2016qzd . These data might help elucidate whether KK production is at the center of the cosmic-ray muon puzzle.
In conclusion, the dark dimension scenario does not provide a theoretical prediction of why the cosmological constant is so small, but it connects it to other scales, namely the size of the extra dimensions, and for the case of $n=2$ to TeV scale physics. Actually, the 6D Planck scale in the $10~{}{\rm TeV}$ regime is on the one hand already quite high in order for being a “natural” solution of the electroweak hierarchy problem and would already need some fine tuning. On the other hand, it is not yet excluded by LHC data, however the next high luminosity LHC run could possibly exclude it. As shown in our paper, the more severe tension of the two extra dark dimension scenario comes from cosmology. A further argument in favor of only one dark dimension comes from gauge unification, which apparently needs a high species scale Heckman:2024trz . Yet another argument favoring $n=1$ comes from axion physics. On the one hand, for a single dark dimension, the QCD axion can be naturally localized on the SM brane, where observational constraints force a narrow range for the axion decay constant $10^{9}\lesssim f_{a}/{\rm GeV}\lesssim 10^{10}$ while the axion mass, $1\lesssim m_{a}/{\rm meV}\lesssim 10$, is surprisingly near the mass scale for the dark energy, the dark matter tower, and the neutrino Gendler:2024gdo . On the other hand, for two dark dimensions of micron size, a large value of $f_{a}$ consistent with observations Hamaguchi:2018oqw could come from the suppression of the wave function if the axion lives in the bulk, but for $T_{*}\sim 2.15~{}{\rm MeV}$, such an axion would be severely constrained through its contribution to $N_{\rm eff}$.
@@ Acknowledgements
We like to thank Allen Caldwell, Ivano Basile, Gia Dvali, Georges Obied, Georg Raffelt, and Cumrun Vafa for useful discussions. We thank Kevork Abazajian for permission to reproduce Fig. 1. The work of L.A.A. is supported by the U.S. National Science Foundation (NSF Grant PHY-2412679). I.A. is supported by the Second Century Fund (C2F), Chulalongkorn University. The work of D.L. is supported by the Origins Excellence Cluster and by the German-Israel-Project (DIP) on Holography and the Swampland. L.A.A. thanks the Theoretical Physics Laboratory at the UNLP in Argentina for hospitality during completion of this work.